# ExPLoRA: Parameter-Efficient Extended Pre-Training to Adapt Vision Transformers under Domain Shifts

### v3

@@ 6.1.2 Ablation study
[L] 0 [] 12.7M 74.83
[L-1,L] 0 [] 25.3M 75.97
[] 256 [Q,V] 25.9M 75.51
[] 128 All 33.1M 55.03
[L] 64 Mlp 16.5M 48.55
[1] 64 [Q,V] 18.7M 75.97
[9] 64 [Q,V] 18.7M 75.45
[L-1] 64 [Q,V] 18.7M 77.40
[L] 0 VPT-100 12.8M 70.14
[L] 64 [Q,V] 18.6M 76.78
[L] 8 [Q,V] 13.4M 76.31
[L] 32 [Q,V] 15.7M 76.40
[L] 64 [Q,V] 18.7M 77.48
Table 3: Ablation study using DinoV2-ExPLoRA, measuring linear-probing accuracy on fMoW-RGB. All results are obtained by using concatenated features from the last 4 ViT blocks.
@@
Next, we investigate whether high LoRA ranks used on all ViT layers (i.e. all attention and MLP matrices, not just $Q,V$) is beneficial. Surprisingly, this significantly harms learning (row 4, 5). In fact, it is much less effective than using just LoRA-$r256$ on the $Q,V$ matrices of all ${\mathcal{L}}$ blocks (row 3). However, both rows 3 and 4 are much less parameter-efficient than ExPLoRA (rows 6-8, 11-13).
@@
Lastly, we investigate the impact of LoRA rank on ExPLoRA. Changing the rank from 8 to 32 has a small improvement ($\uparrow 0.09\%$), but changing from 32 to 64 brings about a much larger improvement ($\uparrow 1.08\%$), with only a relatively small increase in trainable parameters. This demonstrates that higher ranks are necessary during pre-training for effective learning on the new domain. Further ablations on compute efficiency (B.2), data efficiency (B.3), MAE decoder rank (B.4), and ViT backbone size (B.5) are in appendix B.
@@ 6.1.3 Multi-Spectral Satellite Images
Model PEFT
Pre-train/Fine-tune
#Params
@@
MAE [26] Full - / 303.3M 51.61
SatMAE [17] Full 303.3M / 303.3M 61.48
MAE [26] LoRA-$r8$ / 0.8M 46.97
SatMAE [17] LoRA-$r8$ 303.3M / 0.8M 59.48
GFM [44] LoRA-$r8$ 303.3M / 0.8M 57.55
GDA [56] GDA-$r16$ 7.3M / 7.3M 55.23
@@
GASSL [2] Full 74.11
SatMAE [17] Full 79.69
MAE [26] LoRA-r8 69.30
SatMAE [17] LoRA-r8 75.27
M-[L]-$r32$ LoRA-r8 75.98
Table 5: Classification results on the validation set of fMoW-Temporal. All MAE/SatMAE experiments use ViT-L, while GASSL uses a ResNet.
@@
SatMAE [17] Full 78.07 94.80
ScaleMAE [54] Full 78.90 95.70
DinoV2 [47] LoRA-r8 76.69 97.60
D-[L]-$r64$ LoRA-r8 76.69 97.65
SatMAE [17] Lin. Probe 50.89 88.30
ScaleMAE [54] Lin. Probe 47.17 89.60
DinoV2 [47] Lin. Probe 76.21 96.34
D-[L]-$r64$ Lin. Probe 76.34 97.32
Table 6: Results on the validation sets of SpaceNet-v1 (SN), a segmentation task, and Resisc-45 (R45), a classification task.
@@
CLater [53] Full 93.90
ICON Full 90.10
DinoV2 [47] Lin. Probe 93.27
DinoV2 [47] LoRA-r8 92.97
D-[L]-$r32$ Lin. Probe 94.41
D-[L]-$r32$ LoRA-r8 94.21
Table 7: Classification results on the validation set of Camelyon17.
@@
DinoV2 [47] Lin. Probe 66.04
DinoV2 [47] LoRA-r8 67.10
D-[L]-$r32$ Lin. Probe 62.95
D-[L]-$r32$ LoRA-r8 68.07
@@
ICON [34] 68.9 - -
MAE [26] 82.5 53.8 58.7
DinoV2 [47] 82.3 52.1 57.1
D-[L]-$r64$ 82.7 54.5 59.2
Table 9: Object detection results on the validation set of GlobalWheat. AP and AR stand for average precision and average recall.
Dataset
With fMoW-Sentinel, we evaluate transfer from natural images to multi-spectral, low-resolution satellite images - a harder task than fMoW-RGB due to the absence of non-RGB bands in $S$. We use the group-channel ViT-L from Cong et al. (2022), initialized with MAE. During algorithm 1, we additionally unfreeze only the patch embedding layers due to architectural differences.
Table 6 shows the challenge: fully fine-tuning from MAE drops accuracy by nearly 10% (row 1 vs row 2), LoRA tuning from MAE performs worse (row 4), and unfreezing four transformer blocks (row 6) fails to help. However, ExPLoRA with ${\mathcal{U}}=\{\tt{1,L}\}$ outperforms even full pre-training from scratch for LoRA fine-tuning (row 5 vs. last row), demonstrating effective adaptation to a very different domain while using <10% of the parameters.
6.1.4 Additional Satellite Datasets
fMoW-Temporal
Each input is a sequence of up to 3 fMoW-RGB (Christie et al., 2018) images of a location, distributed temporally, and paired with one of 62 classes. Since the inputs are now temporal sequences, we initialize the temporal MAE architecture from Cong et al. (2022) with MAE weights, and pre-train on ${\mathcal{X}}_{T}$ with ${\mathcal{U}}=\tt[L]$ and LoRA rank 32. ExPLoRA then outperforms temporal SatMAE for PEFT (table 6), demonstrating successful transfer learning at a fraction of the pre-training parameters.
SpaceNet-v1
RESISC-45
6.2 WiLDS Datasets
@@
iWildcam classification requires identifying one of 182 animal species given an image. We pre-train on the training set, finding that this outperforms pre-training on the extra-unlabeled set. In table 9, we find an improvement over DinoV2 using LoRA-$r8$ PEFT. Surprisingly, the linear probing performance of the ExPLoRA suffers in comparison with DinoV2, suggesting possible loss in knowledge-transfer due to a small domain gap. Likely because natural image datasets $W_{S}$ such as ImageNet (Deng et al., 2009) used for DinoV2 already contain many images of animals.
@@ 7 Further Analysis
We conduct an investigation on 5 models using a sample of ${\mathcal{X}}_{D_{T}}$. These models are DinoV2, D-ExPLoRA-[L]-$r64$, SatMAE, MAE, and M-ExPLoRA-[L]-$r64$. We do the following analyses: (i) PCA to measure the mean and variance of eigenvalues of patch feature vectors for each ViT block, in figs. 3 and 4 (ii) linear probing for local or global information (Darcet et al., 2023) by training logistic regression classifiers on each block’s patch feature vectors, to predict either patch position (fig. 5) or image class (fig. 6).
@@ 8 Conclusion and Discussion
In this paper, we introduce ExPLoRA, a novel pre-training strategy to adapt pre-trained ViT foundation models for natural images to additional visual domains such as satellite imagery or medical data. We challenge the common paradigm of expensive pre-training from scratch for each new visual domain by offering a solution to transfer knowledge from foundation models that is both parameter-efficient and effective (even outperforming domain-specific foundation models). Our hope is that ExPLoRA enables further use of VFMs on domains other than natural images without requiring vast computational resources for pre-training.
@@ A.1 Comparison with Geospatial Domain Adaptation
We also demonstrate ExPLoRA’s broader applicability through experiments on larger datasets (fMoW-RGB, fMoW-Sentinel, which have 400k-800k images vs 90k images in FireRisk (Shen et al., 2023), the largest dataset used in GDA) and domains beyond remote sensing (i.e. WiLDS). Our analysis in section 7 provides insights into block-wise information encoding, offering practitioners a systematic approach for block selection during PEFT– a unique feature not present in prior work.
@@ Appendix B Additional Experimental Results
We include further experimental results as a continuation of section 6.
@@ B.1 Results on Additional Downstream Datasets
GASSL [2] Full 57.63
SatMAE [17] Full 71.77
SatMAE [17] LoRA-r8 69.45
MAE [26] LoRA-r8 70.36
DinoV2 [47] LoRA-r8 70.40
D-[L]-$r32$ LoRA-r8 70.40
@@
SeCo [43] Full 93.14
SatMAE [17] Full 98.98
SatMAE [17] LoRA-r8 98.73
DinoV2 [47] BOFT-b8m2 96.60
M-[1,L]-$r64$ LoRA-r8 98.54
@@
The dataset contains 27,000 13-band satellite images of 10 classes (Helber et al., 2019), sourced from Sentinel-2. For ExPLoRA, we don’t pre-train on the training set of this dataset, and instead use LoRA fine-tuning starting with the pre-trained weights learned in row 8 of table 6. We demonstrate improved performance over DinoV2, and match the performance achieved by the domain-specific SatMAE which was fully pre-trained on fMoW-Sentinel, and fully fine-tuned on EuroSAT (table 13). This demonstrates the successful use of our extended pre-trained model on further downstream datasets.
@@ B.2 The Importance of Extended Pre-training
Results in fig. 7 demonstrate that ExPLoRA’s extended pre-training achieves a $\uparrow 1.0\%$ improvement in maximum top-1 fine-tuning accuracy within the same total compute budget (320 GPU hours). Notably, even increasing the parameter budget during fine-tuning fails to match this performance. While additional pre-training iterations beyond 50k improve initial fine-tuning accuracy, they have minimal impact on the final accuracy ceiling, highlighting ExPLoRA’s computational efficiency.
@@ B.5 Impact of ViT backbone size
DinoV2 [47] ViT-B 63.62/65.90
DinoV2 [47] ViT-L 67.60/69.00
DinoV2 [47] ViT-G 70.07/70.36
D-[12]-$r64$ ViT-B 74.72/75.11
D-[24]-$r64$ ViT-L 76.86/77.48
D-[32]-$r32$ ViT-G 77.29/77.79
@@ B.6 Additional PEFT baselines for MAE
#Params
@@
SatMAE [17] ViT-L LoRA-$r8$ [28] 303.3M 0.8M 76.10
MAE [26] ViT-L LoRA-$r8$ [28] - 0.8M 76.21
MAE [26] ViT-L DVPT-10 [29] - 0.4M 72.35
MAE [26] ViT-L GVPT-100 [70] - 0.4M 70.86
MAE [26] ViT-L SA2VP [49] - 1.1M 73.55
MAE [26] ViT-L AdaLoRA-$r8$ [72] - 1.2M 75.25
MAE [26] ViT-L Adapter+ [59] - 1.4M 74.10
MAE [26] ViT-L Mona [69] - 7.1M 74.76
M-[L]-$r64$ ViT-L LoRA-$r8$ [28] 18.7M 0.8M 76.55
Table 16: MAE+PEFT results on fMoW-RGB validation split (table 1, contd.). “Pre-train #Params" and “Fine-tune #Params" refer to trainable parameters of the ViT encoder required on the new domain (satellite images).

### v4

@@ 6.1.2 Ablation study
1   – 69.00
2
All
N/A Yes [] 303.3M 1200 54.29
3 [L] 0 [] 12.7M 90 74.83
4 [L-1,L] 0 [] 25.3M 130 75.97
5 [] 256 [Q,V] 25.9M 180 75.51
6 [] 128 All 33.1M 220 55.03
7 [L] 64 Mlp 16.5M 140 48.55
8 [1] 64 [Q,V] 18.7M 100 75.97
9 [9] 64 [Q,V] 18.7M 100 75.45
10 [L-1] 64 [Q,V] 18.7M 100 77.40
11 [L] 0 VPT-100 12.8M 430 70.14
12 [L] 64 [Q,V] 18.6M 100 76.78
13 [L] 8 [Q,V] 13.4M 90 76.31
14 [L] 32 [Q,V] 15.7M 100 76.40
15 [L] 64 [Q,V] 18.7M 100 77.48
16 [L-1,L] 64 [Q,V] 31.1M 140 77.76
17 [1,L-1,L] 64 [Q,V] 43.4M 180 78.04
Table 3: Ablation study using DinoV2-ExPLoRA, measuring linear-probing accuracy on fMoW-RGB. The second row performs full pre-training from scratch. All results are obtained by using concatenated features from the last 4 ViT blocks.
@@
Next, we investigate whether applying high-rank LoRA to all matrices (including MLP) outperforms targeting only attention $Q,V$ matrices. Surprisingly, LoRA-$r128$ on all matrices (row 6) or only on MLP matrices (row 7) significantly harms performance compared to LoRA-$r256$ on just Q,V matrices (row 5). However, both rows 5 and 6 are much less parameter-efficient than ExPLoRA (rows 13-15).
@@
[…]
These results indicate that strategically constraining the parameter space via $\Theta({\mathcal{U}},r)$ in ExPLoRA yields better performance more efficiently than either uniform high-rank updates across all layers (row 5) or simply unfreezing more blocks without LoRA (row 4). This balance between full-rank updates in key transformer blocks and low-rank updates elsewhere more effectively captures domain-specific knowledge during pre-training.
Notably, full pre-training from scratch (row 2) achieves only 54.29% accuracy despite using 1200 GPU hours, suggesting that significantly more compute might be needed for a performant from-scratch DinoV2 checkpoint. In contrast, our standard configuration for ExPLoRA in row 15 achieves 77.48% accuracy with just 100 GPU hours (a 12$\times$ reduction in compute) while using only 18.7M parameters (6% of the full model). While rows 16-17 show further performance improvements with additional unfrozen blocks, row 15 represents our preferred trade-off between performance and efficiency for most applications. Further ablations on compute efficiency (B.2), data efficiency (B.3), MAE decoder rank (B.4), and ViT backbone size (B.5) are in appendix B.
@@ 6.1.3 Multi-Spectral Satellite Images
Dataset
Method PEFT
PT / FT
Params
PT / FT
GPU hours
@@
1    MAE [28] Full – / 303.3M – / 770 51.61
2    SatMAE [17] Full 303.3M / 303.3M 1150 / 770 61.48
3    MAE [28] LoRA-$r8$ – / 0.8M – / 290 46.97
4    SatMAE [17] LoRA-$r8$ 303.3M / 0.8M 1150 / 290 59.48
5    GFM [45] LoRA-$r8$ 303.3M / 0.8M 960 / 290 57.55
6    GDA [58] GDA-$r16$ 7.3M / 7.3M 560 / 410 55.23
@@
With fMoW-Sentinel, we evaluate transfer from natural images to multi-spectral, low-resolution satellite images- a domain with significant distribution shift from RGB images due to the absence of non-RGB bands in $S$. We use the group-channel ViT-L from Cong et al. (2022), initialized with MAE. During ExPLoRA, we additionally unfreeze only the patch embedding layers due to architectural differences.
Table 4 highlights the challenge: direct fine-tuning of MAE weights on this domain results in a substantial 9% performance gap compared to SatMAE (rows 1 vs 2). LoRA tuning from MAE performs worse (row 3), and unfreezing four transformer blocks (row 7) fails to help. ExPLoRA bridges this gap effectively while requiring just 320 GPU hours and <10% trainable parameters for pre-training compared to 1150 hours for SatMAE. This demonstrates ExPLoRA’s ability to efficiently adapt to domains with substantial distribution shifts from natural images, preserving performance while dramatically reducing computational requirements.
6.1.4 Additional Satellite Datasets
@@
1    GASSL [2] Full 23.5M / 23.5M 380 / 220 74.11
2    SatMAE [17] Full 303.3M / 303.3M 1120 / 610 79.69
3    MAE [28] LoRA-$r8$ – / 0.8M – / 250 69.30
4    SatMAE [17] LoRA-$r8$ 303.3M / 0.8M 1120 / 250 75.27
5    M-[L]-$r32$ LoRA-$r8$ 18.7M / 0.8M 100 / 250 75.98
Table 5: Classification results on the validation set of fMoW-Temporal. “PT” and “FT” refer to pre-training and fine-tuning, respectively. All MAE/SatMAE experiments use ViT-L, while GASSL uses a ResNet. ExPLoRA (M-[L]-$r32$) achieves the best PEFT performance while using only a fraction of the compute required for full pre-training.
@@
1        SatMAE [17] Full 78.07 94.80
2        ScaleMAE [56] Full 78.90 95.70
3        DinoV2 [48] LoRA-$r8$ 76.69 97.60
4        D-[L]-$r64$ LoRA-$r8$ 76.69 97.65
5        SatMAE [17] Lin. Probe 50.89 88.30
6        ScaleMAE [56] Lin. Probe 47.17 89.60
7        DinoV2 [48] Lin. Probe 76.21 96.34
8        D-[L]-$r64$ Lin. Probe 76.34 97.32
Table 6: Results on the validation sets of SpaceNet-v1, a segmentation task, and Resisc-45, a classification task. The D-[L]-$r64$ ExPLoRA model from the final row of Table 1 achieves state-of-the-art performance on Resisc-45 and competitive results on SpaceNet-v1 without requiring pre-training.
fMoW-Temporal
Each input is a sequence of up to 3 fMoW-RGB (Christie et al., 2018) images of a location, distributed temporally, and paired with one of 62 classes. Since the inputs are now temporal sequences, we initialize the temporal MAE architecture from Cong et al. (2022) with MAE weights, and pre-train on ${\mathcal{X}}_{T}$ with ${\mathcal{U}}=\tt[L]$ and LoRA rank 32. ExPLoRA then outperforms temporal SatMAE for PEFT (table 5), demonstrating successful transfer learning at a fraction of the pre-training parameters and compute.
SpaceNet-v1
RESISC-45
6.2 WiLDS Datasets
@@
1        CLater [54] Full 93.90
2        ICON Full 90.10
3        MAE [28] LoRA-$r8$ – / 120 92.13
4        M-[L]-$r32$ LoRA-$r8$ 90 / 120 92.24
5        DinoV2 [48] Lin. Probe – / 10 93.27
6        DinoV2 [48] LoRA-$r8$ – / 120 92.97
7        D-[L]-$r32$ Lin. Probe 90 / 10 94.41
8        D-[L]-$r32$ LoRA-$r8$ 90 / 120 94.21
Table 7: Classification results on the validation set of Camelyon17. ExPLoRA models achieve the best performance in both linear probing and LoRA-based fine-tuning configurations.
@@
1        MAE [28] LoRA-$r8$ – / 90 60.07
2        M-[L]-$r32$ LoRA-$r8$ 70 / 90 61.86
3        DinoV2 [48] Lin. Probe – / 10 66.04
4        DinoV2 [48] LoRA-$r8$ – / 90 67.10
5        D-[L]-$r32$ Lin. Probe 70 / 10 62.95
6        D-[L]-$r32$ LoRA-$r8$ 70 / 90 68.07
@@
1        ICON [35] 68.9
2        MAE [28] – / 190 82.5 53.8 58.7
3        M-[L]-$r64$ 30 / 190 79.3 51.6 56.6
4        DinoV2 [48] – / 190 82.3 52.1 57.1
5        D-[L]-$r64$ 30 / 190 82.7 54.5 59.2
Table 9: Object detection results on the validation set of GlobalWheat. AP and AR stand for average precision and average recall. ExPLoRA outperforms the baseline methods across all metrics.
@@
iWildcam classification requires identifying one of 182 animal species given an image. We pre-train on the training set, finding that this outperforms pre-training on the extra-unlabeled set. In table 8, we find an improvement over DinoV2 using LoRA-$r8$ PEFT. Surprisingly, the linear probing performance of the ExPLoRA suffers in comparison with DinoV2, suggesting possible overfitting due to a small domain gap. This is likely because natural image datasets such as ImageNet (Deng et al., 2009) used for pre-training DinoV2 already contain many images of animals.
@@ 7 Further Analysis
We conduct an investigation on 5 models using a sample of ${\mathcal{X}}_{D_{T}}$. These models are DinoV2, D-ExPLoRA-[L]-$r64$, SatMAE, MAE, and M-ExPLoRA-[L]-$r64$. We do the following analyses: (i) PCA to measure the mean and variance of eigenvalues of patch feature vectors for each ViT block, in figs. 3 and 4 (ii) linear probing for local or global information (Darcet et al., 2024) by training logistic regression classifiers on each block’s patch feature vectors, to predict either patch position (fig. 5) or image class (fig. 6).
@@
Thus, our analysis provides guidelines for practitioners to select which blocks to unfreeze based on the eigenvalue properties and classification accuracy patterns of different ViT layers, offering a systematic approach to constraining $\Theta({\mathcal{U}},r)$ when pre-training with ExPLoRA (eq. 5).
@@ 8 Conclusion and Discussion
In this paper, we introduce ExPLoRA, a novel pre-training strategy to adapt pre-trained ViT foundation models for natural images to additional visual domains such as satellite imagery or medical data. We challenge the common paradigm of expensive pre-training from scratch for each new visual domain by offering a solution to transfer knowledge from foundation models that matches or outperforms domain-specific foundation models. ExPLoRA makes powerful foundation models accessible to researchers with limited computational resources while using 8-10x less compute and 16x fewer parameters. Our hope is that ExPLoRA enables further use of VFMs on domains other than natural images without requiring vast computational resources for pre-training.
@@ A.1 Comparison with Geospatial Domain Adaptation
We also demonstrate ExPLoRA’s broader applicability through experiments on larger datasets (fMoW-RGB, fMoW-Sentinel, which have 400k-800k images vs 90k images in FireRisk (Shen et al., 2023), the largest dataset used in GDA) and domains beyond remote sensing (i.e. WiLDS). Our analysis in section 7 provides insights into block-wise information encoding, offering practitioners a systematic approach for block selection during extended pre-training– a unique feature not present in prior work.
@@ Appendix B Additional Experimental Results
Below, we include further experimental results as a continuation of section 6.
@@ B.1 Results on Additional Downstream Datasets
1         GASSL [2] Full 57.63
2         SatMAE [17] Full 71.77
3         SatMAE [17] LoRA-r8 69.45
4         MAE [28] LoRA-r8 70.36
5         DinoV2 [48] LoRA-r8 70.40
6         D-[L]-$r32$ LoRA-r8 70.40
@@
1        SeCo [44] Full 93.14
2        SatMAE [17] Full 98.98
3        SatMAE [17] LoRA-r8 98.73
4        DinoV2 [48] BOFT-b8m2 96.60
5        M-[1,L]-$r64$ LoRA-r8 98.54
@@
The dataset contains 27,000 13-band satellite images of 10 classes (Helber et al., 2019), sourced from Sentinel-2. For ExPLoRA, we don’t pre-train on this dataset’s training set, and instead use LoRA fine-tuning starting with the pre-trained weights learned in the last row of table 4. We demonstrate improved performance over DinoV2, and match the performance achieved by the domain-specific SatMAE which was fully pre-trained on fMoW-Sentinel, and fully fine-tuned on EuroSAT (table 13). This demonstrates the successful use of our extended pre-trained model on further downstream datasets.
@@ B.2 The Importance of Extended Pre-training
Results in fig. 7 demonstrate that ExPLoRA’s extended pre-training achieves a $\uparrow 1.0\%$ improvement in maximum fine-tuning accuracy within the same total compute budget (320 GPU hours). Notably, even increasing the parameter budget during fine-tuning fails to match this performance. While additional pre-training iterations beyond 100k improve initial fine-tuning accuracy, they have minimal impact on the final accuracy ceiling, highlighting ExPLoRA’s computational efficiency. Importantly, the total amount of compute invested in pre-training and fine-tuning ExPLoRA is still dwarfed by the compute spent for full pre-training from scratch, e.g., 960 GPU hours for SatMAE (table 1).
Lastly, we re-iterate that the benefits of extended pre-training become more pronounced as the domain gap increases. While fig. 7 demonstrates a 1% improvement on RGB satellite images, table 4 highlights a 8% improvement when using ExPLoRA + PEFT (row 8) over full fine-tuning MAE (row 1) on multi-spectral satellite images– a domain with substantial shift from natural images. ExPLoRA still requires less total compute (610 vs 770 GPU hours), demonstrating that its efficiency advantages scale with domain difficulty.
@@ B.5 Impact of ViT backbone size
1      DinoV2 [48] ViT-B 63.62/65.90
2      DinoV2 [48] ViT-L 67.60/69.00
3      DinoV2 [48] ViT-G 70.07/70.36
4      D-[12]-$r64$ ViT-B 74.72/75.11
5      D-[24]-$r64$ ViT-L 76.86/77.48
6      D-[32]-$r32$ ViT-G 77.29/77.79
@@ B.6 Additional PEFT baselines for MAE
Params
Pre-Train
GPU hours
Fine-Tune
GPU hours
@@
1         SatMAE [17] ViT-L LoRA-$r8$ [30] 303.3M 0.8M 960 220 76.10
2         MAE [28] ViT-L LoRA-$r8$ [30] 0.8M 220 76.21
3         MAE [28] ViT-L DVPT-10 [31] 0.4M 500 72.35
4         MAE [28] ViT-L GVPT-100 [74] 0.4M 500 70.86
5         MAE [28] ViT-L SA2VP [50] 1.1M 410 73.55
6         MAE [28] ViT-L AdaLoRA-$r8$ [77] 1.2M 220 75.25
7         MAE [28] ViT-L Adapter+ [61] 1.4M 220 74.10
8         MAE [28] ViT-L Mona [73] 7.1M 220 74.76
9         M-[L]-$r64$ ViT-L LoRA-$r8$ [30] 18.7M 0.8M 80 220 76.55
Table 16: MAE+PEFT results on fMoW-RGB validation split (table 1, contd.). “Pre-train Params" and “Fine-tune Params" refer to trainable parameters of the ViT encoder required on the new domain (satellite images).
Pre-training ExPLoRA with MAE is slightly more efficient than with DinoV2 (i.e. 80 vs 100 GPU hours) due to two factors: the MAE encoder only operates on visible image patches, and unlike DinoV2, it does not require maintaining a separate "teacher" model copy.